Calculate the annual dividend growth rate for the last 10 years. Based on the annual dividend growth rate for the company, will you forecast a constant or non-constant growth in dividends? Explain.

ROP Company

Date Dividends
1/6/10 0.095
4/7/10 0.095
7/7/10 0.095
10/6/10 0.095
1/5/11 0.11
4/6/11 0.11
7/6/11 0.11
10/5/11 0.11
1/4/12 0.1375
4/11/12 0.1375
7/11/12 0.1375
10/10/12 0.1375
12/19/12 0.165
4/10/13 0.165
7/10/13 0.165
10/9/13 0.165
1/8/14 0.2
4/9/14 0.2
7/9/14 0.2
10/8/14 0.2
1/7/15 0.25
4/8/15 0.25
7/8/15 0.25
10/7/15 0.25
1/6/16 0.3
4/6/16 0.3
7/6/16 0.3
10/5/16 0.3
1/5/17 0.35
4/5/17 0.35
7/5/17 0.35
10/5/17 0.35
1/8/18 0.4125
4/6/18 0.4125
7/6/18 0.4125
10/4/18 0.4125
1/8/19 0.4625
Answer #1

Solution:

Annual dividend growth rate for last 10 years.

Based on the information provided:

0.095 as on 01/06/10

0.4625 as on 1/8/19

0.4625 / 0.095 = 4.4684

\10\sqrt[]{4.4684} = 21.1385

21.1385 - 1 = 20.1385

20.1385 * 100 = 2013.85 % (Answer) --------Annual dividend growth rate for last 10 years

There has been a significanr divident growth in dividends, has increased over a period of time and then remained constant for few quarters, then again increased and so on.

Based on historical data--------company will incease dividend in both ways.